OpenStudy (anonymous):

does the table represent an exponential function? x 1 2 3 4 y 4 16 64 256

OpenStudy (godlovesme):

if u look closely do u see a pattern? @Alice2112

OpenStudy (godlovesme):

\[4*4=16\] \[16*4=64\] \[64*4=256\]

OpenStudy (godlovesme):

so it's multiplying each # by 4 confused? @Alice2112

OpenStudy (anonymous):

okay i understand

OpenStudy (godlovesme):

okiee so is it exponential?

OpenStudy (anonymous):

i want to say yes but im not sure what exponential is @Godlovesme

OpenStudy (godlovesme):

exponential is if the pattern is multiplication like the one above